You have a very concentrated solution (12 M) of potassium chloride (KCl). You need it to be at the least concentration possible for the experiment you are about to conduct. The problem is you forgot to order lab supplies, so you only have 2 L of distilled water left. What would be the final concentration if you added the two 2 L of distilled water to the 0.5 L of 12 M KCl?
1
Expert's answer
2019-03-06T07:46:28-0500

Cm= n/V

Cm1=12M

Cm2=?

n=Cm x V= 12 x 0.5 = 6 mol

Cm2= 6 /2.5=2.4 M